OpenStudy (anonymous):

write this equation in standard form. -2+x squared=-x. show work please :)

OpenStudy (anonymous):

Change them to standard form which is Ax + By = C where A and B are not zero AND A is a positive number: -2+x^2=-x +x +x ------------------- -2+x^2+x= +2 +2 --------------------- x^2+x=2 I think that is correct